For what values of k does the line y = kx –4 pass through the point of intersection of the lines y=2x–5 and y=–x+1?
Found 2 solutions by mananth, robertb:
Answer by mananth(16946)

2.00 x -1.00 y = 5.00 .............1
Total value
1.00 x + 1.00 y = 1.00 .............2
Eliminate y
multiply (1)by 1.00
Multiply (2) by 1.00
2.00 x -1.00 y = 5.00
1.00 x + 1.00 y = 1.00
Add the two equations
3.00 x = 6.00
/ 3.00
x = 2.00
plug value of x in (1)
2.00 x -1.00 y = 5.00
4.00 -1.00 y = 5.00
-1.00 y = 5.00 -4.00
-1.00 y = 1.00
y = -1.00
(2,-1) lies on the line y=kx-4
plug (2,-1)
-1=2k-4
2k=3
k=3/2

Answer by robertb(5830)

The lines y=2x–5 and y=–x+1 intersect at the point (2,-1). (Use method of substitution!)
Substituting the coordinates of (2,-1) into the equation y = kx - 4, we get
-1 = 2k - 4
==> 3 = 2k, or k = 3/2.